A Bibliography of Publications of Edward W. Leaver

Edward W. Leaver Route 52 Fishkill Glen #11G

Fishkill, NY 12524

USA Tel: +1 914 897-5829

FAX: ?n/a?

E-mail: leaver@math.utah.edu (Internet)

25 November 2011 Version 1.06

Abstract

This bibliography records publications of Edward W. Leaver.

Title word cross-reference

analytic [Lea85]. black [Lea85]. center [Lea86a]. Comment [Lea92a]. computing [Lea92b]. continued [Lea92b]. decomposition [Lea86b]. equation [Lea86a]. equations [Lea86a]. fraction [Lea92b]. frequencies [Lea92b]. general [Lea86a]. generalized [Lea86a]. geometry [Lea86b].

High [Lea92a]. High-overtone [Lea92a]. hole [Lea92b]. holes [Lea85]. Kerr [Lea85]. mechanics [Lea86a]. method [Lea92b]. modes [Lea85]. molecular [Lea86a]. Nordström [Lea90]. normal [Lea92a]. overtone [Lea92a]. perturbation [Lea86b]. problem [Lea86a]. quantum [Lea86a]. quasinormal [Lea85]. Reissner [Lea90]. relativity [Lea86a]. Remarks [Lea92b]. representation [Lea85]. response [Lea86b]. Schwarzschild [Lea92a, Lea86b]. Solutions [Lea86a]. Spectral [Lea86b]. spheroidal [Lea86a]. Teukolsky [Lea86a]. two [Lea86a]. two-center [Lea86a].

wave [Lea86a].

References

Leaver:1985:ARQ

[Lea85] Edward W. Leaver. An analytic representation for the quasinormal modes of Kerr black holes. Proc. R. Soc. Lond. A., 402:285–298, 1985.

Leaver:1986:SGS

[Lea86a] Edward W. Leaver. Solutions to a generalized spheroidal wave equation: Teukolsky’s equations in general relativity, and the two-center problem in molecular quantum mechanics. J. Math and Physics, 27: 1238–1265, 1986.

Leaver:1986:SDP

[Lea86b] Edward W. Leaver. Spectral decomposition of the perturbation response of the Schwarzschild geometry. Phys. Rev. D, 34:384–408, 1986.

Leaver:1990:QMR

[Lea90] Edward W. Leaver. Quasinormal modes of Reissner–Nordström black holes. Phys. Rev. D, 41:2986, 1990.

Leaver:1992:CON

[Lea92a] Edward W. Leaver. Comment on “High-overtone normal modes of Schwarzschild black holes”. Class. Quantum Grav., 9:1643–1648, 1992.

Leaver:1992:RCF

[Lea92b] Edward W. Leaver. Remarks on the continued fraction method for computing black hole quasinormal frequencies and modes. Phys. Rev. D, 45:4713, 1992.